A vertical line passing through the point (1,2) intersects the X axis at A(a, 0) and Y axis at B(0, b). Find area of triangle of least area if a and b are positive.


1
Expert's answer
2021-11-08T08:12:00-0500

Equation of AB ( intercept form)


"\dfrac{x}{a}+\dfrac{y}{b}=1"

A line passes through the point "(1,2)"


"\dfrac{1}{a}+\dfrac{2}{b}=1"

"\dfrac{1}{a}=1-\dfrac{2}{b}"

"a=\dfrac{b}{b-2}, a>0, b>2"

The area of "\u25b3AOB" is


"Area=\dfrac{ab}{2}"

"Area=A(b)=\dfrac{b^2}{2(b-2)}"

Find the first derivative with respect to "b"


"A'(b)=(\dfrac{b^2}{2(b-2)})'=\dfrac{2b(b-2)-b^2}{2(b-2)^2}=\dfrac{b^2-4b}{2(b-2)^2}"

Find the critical number(s)


"A'(b)=0=>\dfrac{b^2-4b}{2(b-2)^2}=0=>b_1=0, b_2=4"

We consider "b>2."

If "2<b<4," then "A'(b)<0,A(b)" decreases.


If "b>4," then "A'(b)>0,A(b)" increases.

The function "A(b)" has a local minimum at "b=4."

Since the function "A(b)" has the only extremum for "b>2," then the function "A(b)" has the absolute minimum for "b>2" at "b=4."


"a=\dfrac{4}{4-2}=2"

"Area_{min}=\dfrac{2(2)}{2}=2(square\ units)"